# TopoBenchmarkX: A Framework for Benchmarking Topological Deep Learning

### v1

@@ Abstract
This work introduces TopoBenchmarkX, a modular open-source library designed to standardize benchmarking and accelerate research in Topological Deep Learning (TDL). TopoBenchmarkX maps the TDL pipeline into a sequence of independent and modular components for data loading and processing, as well as model training, optimization, and evaluation. This modular organization provides flexibility for modifications and facilitates the adaptation and optimization of various TDL pipelines. A key feature of TopoBenchmarkX  is that it allows for the transformation and lifting between topological domains. This enables, for example, to obtain richer data representations and more fine-grained analyses by mapping the topology and features of a graph to higher-order topological domains such as simplicial and cell complexes. The range of applicability of TopoBenchmarkX is demonstrated by benchmarking several TDL architectures for various tasks and datasets.
@@ 1 Introduction
In geometric deep learning (GDL; Bronstein et al., 2021), graph neural networks (GNNs; Zhou et al., 2020) have demonstrated impressive capabilities in processing relational data represented as graphs. However, graphs only encode pairwise structures, which can be limiting. For example, social interactions may involve multiple individuals, and electrostatic interactions in proteins may involve multiple atoms. Topological deep learning (TDL; Papamarkou et al., 2024; Bodnar, 2023; Hajij et al., 2023; Papillon et al., 2023b) offers a framework for modeling complex systems characterized by such multi-way (higher-order) relations among components. Topological neural networks (TNNs; Feng et al., 2019; Bunch et al., 2020; Hajij et al., 2020; Bodnar et al., 2021a; Ebli et al., 2020; Schaub et al., 2021; Bodnar et al., 2021b; Chien et al., 2021), which are part of TDL, have found applications in numerous fields that involve higher-order relational data such as social networks (Knoke and Yang, 2019) and protein biology (Jha et al., 2022). TNNs achieve state-of-the-art performance in various machine learning tasks (Dong et al., 2020; Barbarossa and Sardellitti, 2020; Chen et al., 2022; Roddenberry et al., 2021; Telyatnikov et al., 2023; Giusti et al., 2023a).
@@
Figure 1: Workflow of TopoBenchmarkX, which consists of four main components, organized by their functionality: data modules, model modules, training modules, and communication modules.
TDL has evolved rapidly recently, providing novel algorithms for learning higher-order relational data supported on various topological domains, such as simplicial, cell, and combinatorial complexes. The works of Hajij et al. (2023) and Papillon et al. (2023b) have recently unified theory and notation of many TNNs. Moreover, the TAG ICML challenge 2023 (Papillon et al., 2023a) has contributed software implementations of topological message passing for most TNNs under TopoX (Hajij et al., 2024). Despite these theoretical and practical advancements, the rapid acceleration of TDL research poses challenges for the reproducibility and systematic comparative evaluation of TNNs (Papamarkou et al., 2024). In particular, open software, standardization, and benchmarking were recognized as key open challenges for TDL in a recent ICML 2024 position paper (Papamarkou et al., 2024). TopoBenchmarkX provides a robust framework to address these issues and ensure consistent and reliable progress in the field.
TopoBenchmarkX [^1] https://github.com/pyt-team/TopoBenchmarkX is an open-source and modular benchmarking framework for comparative evaluations of TNNs. Figure 1 visualizes the overall workflow and modularity of TopoBenchmarkX; see Section 4 for a more detailed outline. TopoBenchmarkX can generate topological datasets for benchmarking purposes, and carry out TNN comparisons based on various performance metrics. TopoBenchmarkX provides a high-level interface that reduces coding efforts, making it user-friendly. It can facilitate reproducibility of TNN research, and serves as a testbed to experiment with new ideas in TDL. Finally, TopoBenchmarkX is part of a larger suite of packages, TopoX (Hajij et al., 2024), and it is built with compatibility in mind, allowing integration with other tools and frameworks both within TopoX and in the broader PyTorch ecosystem.
Three aspects of TDL make benchmarking less straightforward: the scarcity of topological data, the non-trivial standardization of input and output across topological domains, and the diversity of TNN architectures. TopoBenchmarkX addresses these three limitations.
Generation of topological data. Although higher-order data arise in various physical systems, the complexity of capturing these interactions and the limitations of current experimental designs constrain data collection. As a result, higher-order data is often generated by lifting (i.e., transforming) graph datasets and equipping them with higher-order features. TopoBenchmarkX implements graph lifting algorithms to automate this generation process, thus mitigating the scarcity of topological datasets.
Standardization of input and output across topological domains. Each topological domain encodes higher-order relations differently. Further, each TNN architecture provides distinct mappings between inputs and outputs, and therefore processes higher-order data differently. TopoBenchmarkX standardizes the input-output pipeline through an interface that automatically manages the transition between topological domains and their associated TNNs.
@@
Diversity of TNN architectures. There is a broad spectrum of TNNs in the TDL literature. Each TNN typically employs its own techniques to preprocess and encode data. This diversity complicates performance comparisons between models on different datasets. To address this issue, TopoBenchmarkX implements a pipeline for data preprocessing and predictive performance evaluation metrics. The pipeline is easily accessible to newcomers through a configuration file.
The rest of the paper is structured as follows. Section 2 outlines existing software for graph-based learning and benchmarking, as well as GDL. Section 3 introduces the notion of featured cell complex and formalizes the lifting of a graph to a cell complex. Section 4 introduces the main components of TopoBenchmarkX, and Section 5 demonstrates its usage via benchmarking experiments. The paper concludes with some final remarks, current limitations, and suggestions for future work in Section 6.
@@ 2 Existing software
Graph-based learning and GDL are supported by several software packages, including NetworkX (Hagberg et al., 2008), KarateClub (Rozemberczki et al., 2020), PyG (Fey and Lenssen, 2019), and DGL (Wang et al., 2019). NetworkX enables computations on graphs, while KarateClub implements unsupervised learning algorithms for graph-structured data. PyG and DGL provide functionality for GDL as well as graph-based learning.
@@
Various software packages implement machine learning algorithms for higher-order domains such as HyperNetX (Liu et al., 2021), XGI (Landry et al., 2023), DHG (Feng et al., 2019), and TopoX (Hajij et al., 2024). HyperNetX facilitates computing with hypergraphs, while XGI enables computing with both hypergraphs and simplicial complexes. DHG implements deep learning algorithms for graphs and hypergraphs. TopoX is a suite of three packages: TopoNetX, TopoEmbedX, and TopoModelX, which together enable learning with TNNs. TopoNetX supports computations on graphs and higher-order domains, including simplicial complexes, cell complexes, path complexes, combinatorial complexes, and colored hypergraphs. TopoEmbedX implements algorithms to embed higher-order domains into Euclidean spaces. TopoModelX implements most of the TNNs surveyed in Papillon et al. (2023b).
The Open Graph Benchmark (OGB) is a set of datasets and a benchmark framework to facilitate reproducible graph machine learning research (Hu et al., 2020, 2021). The OGB is applicable to graph-based learning and GDL, but not to TDL. TopoBenchmarkX aims to fill this gap by providing algorithms that generate higher-order datasets and benchmark TNNs.
3 Featured topological domains and lifting
This section extends the concept of a featured graph to featured topological domains. To make the new definitions precise, cell complexes are used as a template for other topological domains. However, all definitions provided here are applicable to other complexes implemented in TopoBenchmarkX. This section also formalizes the lifting of a featured graph to a featured topological domain. See Appendix D for an extended introduction to other higher-order domains (e.g., simplicial complexes, hypergraphs).
3.1 Featured topological domains
A featured graph is a graph equipped with functions that map nodes or edges to features Sanchez-Lengeling et al. (2021).
Definition 1
Let $\mathcal{G}=(V,E)$ be a graph, with node set $V$ and edge set $E$. A featured graph is a tuple $\mathcal{G}_{F}=(V,E,F_{V},F_{E})$, where $F_{V}:V\to\mathbb{R}^{d}$ is a function that maps each node to a feature vector in $\mathbb{R}^{d}$ and $F_{E}:E\to\mathbb{R}^{k}$ is a function that maps each edge to a feature vector in $\mathbb{R}^{k}$.
Definition 2
A cell complex is a triplet $\mathcal{X}=(\mathcal{X}_{0},\mathcal{X}_{1},\mathcal{X}_{2})$ of finite ordered sets, where $\mathcal{X}_{0}$ is a set of nodes (or 00-cells), $\mathcal{X}_{1}$ is a set of edges (or $1$-cells) with each edge $e\in\mathcal{X}_{1}$ being an ordered pair of nodes $e=[v_{1},v_{2}]$ for $v_{1},v_{2}\in\mathcal{X}_{0}$, and $\mathcal{X}_{2}$ is a set of faces (or $2$-cells) with each face $\sigma\in\mathcal{X}_{2}$ being an ordered sequence of edges $\sigma=[e_{1},e_{2},\dots,e_{m(\sigma)}]$ forming a closed path without self-intersections and $m(\sigma)\geq 3$. A featured cell complex is a tuple $\mathcal{X}_{F}=(\mathcal{X}_{0},\mathcal{X}_{1},\mathcal{X}_{2},F_{\mathcal{X}_{0}},F_{\mathcal{X}_{1}},F_{\mathcal{X}_{2}}),$ where, for $0\leq i\leq 2$, $F_{\mathcal{X}_{i}}:\mathcal{X}_{i}\to\mathbb{R}^{k_{i}}$ maps each cell in $\mathcal{X}_{i}$ to a feature vector $\mathbb{R}^{k_{i}}$.
3.2 Liftings
In this paper, lifting refers to the process of mapping a featured graph to a featured topological domain through a well-defined procedure (Hajij et al., 2023). We provide a precise definition for the case of cell complexes. However, lifting to other topological domains can be defined similarly.
Definition 3
Let $\mathcal{G}_{F}=(V,E,F_{V},F_{E})$ be a featured graph. A lifting of $\mathcal{G}_{F}$ is a triplet $(\mathcal{C}_{F},\iota,L)$, where $\mathcal{C}_{F}=(\mathcal{X}_{0},\mathcal{X}_{1},\mathcal{X}_{2},F_{\mathcal{X}_{0}},F_{\mathcal{X}_{1}},F_{\mathcal{X}_{2}}),$ is a featured cell complex, $\iota$ is an embedding map that maps the nodes and edges of $\mathcal{G}$ to cells in $\mathcal{X}$, and $L$ is a collection of functions that map features of $\mathcal{G}_{F}$ to features of $\mathcal{C}_{F}$. The embedding map $\iota$ satisfies $\iota(v)=s_{v}$ for $v\in V$ with $s_{v}\in\mathcal{X}_{0}$ being a 00-cell corresponding to node $v$, and $\iota(e)=x_{e}$ for $e\in E$ with $x_{e}\in\mathcal{X}_{1}$ being a cell corresponding to edge $e$. The collection of functions in $L$ consists of a lifting procedure for the higher-order cells in $\mathcal{X}$ that are not part of $\mathcal{G}$ and a lifting procedure for the feature maps $\{F_{\mathcal{X}^{i}}\}_{i=0}^{2}$ based on $\mathcal{G}_{F}$.
Examples of liftings are given in Figure 2. Observe that a featured cell complex in Definition 3 acts as a template for a general topological domain to which the structural and feature information of the graph is lifted. In the general case, even if the graph features are empty (that is, if $F_{V}$ and $F_{E}$ are not defined), lifting still provides a way to map the graph structure to the cell complex structure.
Figure 2: An illustration of lifting a featured graph $\mathcal{G}_{F}$ (center) to two different topological domains: a simplicial complex (left) and a cell complex (right). The lifting from $\mathcal{G}_{F}$ (center) to the featured cell complex $\mathcal{C}_{F}$ (right) is represented by the triplet $(\mathcal{C}_{F},\iota,L)$, where $\iota$ is the embedding map that places the nodes and edges of $\mathcal{G}_{F}$ into the corresponding cells of $\mathcal{C}_{F}$, and $L$ is the collection of functions that maps the features of $\mathcal{G}_{F}$ to the features of $\mathcal{C}_{F}$.
Lifting maps can be fixed (Bodnar et al., 2021a; Hajij et al., 2023) or learnable (Battiloro et al., 2023a; Bernárdez et al., 2023). Fixed lifting maps are predefined and follow a rule-based procedure to lift a featured graph to a featured topological domain. For example, one could define a fixed lifting that always maps a node in the graph to a node in the topological domain, and an edge to a corresponding cell. In contrast, learnable lifting maps are determined using machine learning models that are trained to optimize the lifting process based on specific criteria or objective functions. During training, the model learns to embed the nodes and edges of the input graph into the cells of a topological domain in a way that optimizes the representation for downstream tasks, such as classification or clustering. When lifting a graph to a topological domain, both the features on the higher-order cells and the structure of the topological domain itself can be computed deterministically (Ramamurthy et al., 2023) or learned (Hajij et al., 2023). [^2] An unpooling operation from lower-rank cells, where features are given, to higher-rank cells, where features are to be lifted, is called a learnable feature lifter in Hajij et al. (2023).
4 Outline of TopoBenchmarkX library
@@
TopoBenchmarkX implements a unified and flexible workflow that facilitates the addition of new datasets, transforms, and deep learning models, thus ensuring applicability across a wide range of tasks and enabling a broad cross-domain comparison, which is lacking in the TDL literature. Each module within TopoBenchmarkX is assigned a distinct role, while maintaining a consistent input-output structure, which provides a modular interface across all topological domains. Figure 1 outlines the TopoBenchmarkX modules, which are organized by functionality to data, model, training, and communication modules. Algorithm 1 exemplifies the TopoBenchmarkX execution pipeline via pseudo-code.
4.1 TopoBenchmarkX modules
Data modules. These modules include Loader, Transforms, PreProcessor, and Dataloader.
Loader. The loader module provides an interface for downloading and storing data; it includes the GraphLoader, HypergraphLoader, SimplicialLoader, and CellLoader classes. Their interfaces are built upon the widely-adopted InMemoryDataset, provided by PyG to enhance interoperability. The project webpage provides multiple in-depth tutorials on the library, including a step-by-step overview of incorporating customized data with these interfaces.
Transforms. Transforms modules are developed as sub-classes of the BaseTransform class provided by PyG. These modules are data manipulation, topology lifting, and feature lifting. The data manipulation module allows for performing any data transformation; e.g., the transformations available in PyG or TopoEmbedX can be easily adapted in TopoBenchmarkX. The topology lifting and feature lifting modules are responsible for lifting data from one topological domain to another. In particular, the former module relies on TopoNetX to lift the topology, while the latter utilizes the lifted topology to infer (e.g., project, concatenate) the features for higher-order topological structures. Each transform takes a Data object from PyG as input, performs the necessary computations, and outputs the modified Data object. The Transforms modules are composable operators that can be easily customized to perform various tasks.
Pre-processor. The PreProcessor class applies a sequence of transforms to a dataset. It takes a dataset object and a sequence of transforms as input, and iterates over the dataset, applying a set of transforms sequentially to each instance in the dataset. To avoid re-computing the same transforms multiple times, the pre-processed dataset is automatically saved in a unique folder for each transform configuration. This ensures that each dataset needs to be processed only once per configuration, addressing the potential time-consuming nature of preprocessing large datasets. PreProcessor also generates and/or loads data splits according to the chosen strategy; for example, random splits with predefined proportions, k-fold cross-validation, or fixed splits.
Dataloader. The dataloader module provides a consistent interface for batch training across all topological domains. TopoBenchmarkX supports mini-batching for graphs, hypergraphs, simplicial complexes, and cell complexes.
Model modules. Neural networks modules form the core of the modeling pipeline. The encoder module maps initial data features into a latent space and applies a learnable transformation before passing the data to the TNN model. The backbone TNN can be imported from TopoModelX, PyG, or it can be built on a custom basis, allowing flexibility in model selection (see Table 6 in Appendix E.2). The wrapper ensures that the correct input is provided to the forward pass of the model and collects the output in a dictionary. The wrapper plays a crucial role in the integration of any model in TopoBenchmarkX, making it possible to use existing and newly developed models. The readout module processes latent representations of the neural network into final output predictions. The Loss module is responsible for defining the loss function from the PyTorch library. The optimizer module defines the optimizer and scheduler. It allows for easy integration of any optimizer and scheduler from torch.optim, which provides flexibility in optimizing model training. The evaluator module is responsible for computing metrics during training and inference. It is built upon torchmetrics and supports standard metrics for classification and regression tasks.
Training and communication modules. The Model class defines a training pipeline for all domains (lines 14-19 of Algorithm 1). It inherits from LightningModule, and requires the Encoder, Wrapper, Backbone, Readout, Evaluator, Loss, and Optimizer objects as inputs. The lightning.Trainer automates the training, evaluation, and testing processes. Extra functionalities can be added using callbacks modules, and users can monitor the training process using different loggers (e.g. wandb, tensorboard). Both are common practice in Lightning, and they are referred to as communication modules in TopoBenchmarkX.
Algorithm 1 Execution pipeline for model training in TopoBenchmarkX

### v2

# TopoBench: A Framework for Benchmarking Topological Deep Learning

@@ Abstract
This work introduces TopoBench, an open-source library designed to standardize benchmarking and accelerate research in topological deep learning (TDL). TopoBench decomposes TDL into a sequence of independent modules for data generation, loading, transforming and processing, as well as model training, optimization and evaluation. This modular organization provides flexibility for modifications and facilitates the adaptation and optimization of various TDL pipelines. A key feature of TopoBench is its support for transformations and lifting across topological domains. Mapping the topology and features of a graph to higher-order topological domains, such as simplicial and cell complexes, enables richer data representations and more fine-grained analyses. The applicability of TopoBench is demonstrated by benchmarking several TDL architectures across diverse tasks and datasets.
Keywords: Benchmark, topological deep learning, topological neural networks.
@@ 1 Introduction
In geometric deep learning (GDL; Bronstein et al., 2021), graph neural networks (GNNs; Zhou et al., 2020) have demonstrated impressive capabilities in processing relational data represented as graphs. However, because graphs represent relationships through edges, they inherently capture only pairwise interactions, which can be a limiting factor. For example, social interactions often involve groups of individuals rather than just pairs, and electrostatic interactions in proteins can span multiple atoms. Topological deep learning (TDL; Papamarkou et al., 2024; Bodnar, 2023; Hajij et al., 2023; Papillon et al., 2023) offers a framework for modeling complex systems characterized by such multi-way relations among components, leveraging to that end higher-order discrete topological domains (such as simplicial and cell complexes, see Section 2). Topological neural networks (TNNs; Feng et al., 2019; Bunch et al., 2020; Hajij et al., 2020; Bodnar et al., 2021a; Ebli et al., 2020; Schaub et al., 2021; Bodnar et al., 2021b; Chien et al., 2021), which are part of TDL, have found applications in numerous fields that involve higher-order relational data such as social networks (Knoke and Yang, 2019), protein biology (Jha et al., 2022), physics (Wei and Fink, 2024), and computer networks (Bernárdez et al., 2025). TNNs have also shown their potential in various machine learning tasks (Dong et al., 2020; Barbarossa and Sardellitti, 2020; Chen et al., 2022; Roddenberry et al., 2021; Telyatnikov et al., 2023; Giusti et al., 2023).
However, as identified in a recent position paper (Papamarkou et al., 2024), the rapid growth of TDL research has introduced challenges in ensuring reproducibility and conducting systematic comparative evaluations of TNNs. To address these challenges, this work introduces TopoBench  [^1] https://github.com/geometric-intelligence/TopoBench, an open-source and modular framework for TDL. By providing a comprehensive pipeline –from data integration and processing to modeling and evaluation–, our proposed framework facilitates both development and benchmarking of TNNs (Figure 1 illustrates the overall workflow). More specifically, TopoBench directly addresses the following relevant limitations of current TDL models’ evaluations (Papamarkou et al., 2024):
@@
Figure 1: Workflow of TopoBench, consisting of four main components: data modules, model modules, training modules, and communication modules.
Data availability: Although many complex systems exhibit higher-order interactions, they are mostly collected in the form of point clouds or graphs, implying the failure to fully capture a more nuanced interplay. For instance, in a social network, we might track friendships between individuals but overlook whether they belong to the same group, losing valuable higher-order relationships. This limitation arises because current experimental designs often impose constraints on what data can be collected, making it difficult to systematically capture complex, multi-level relationships. TopoBench mitigates the scarcity of higher-order data in three ways. First, it provides an interface for uploading publicly available higher-order datasets. Second, it facilitates the loading of user-defined datasets – whether higher-order or not. Third, it implements lifting algorithms (i.e. mappings between different discrete topological domains) to automate the construction of new topological datasets.
@@
Standardization: There is a broad spectrum of TNNs in the TDL literature, each using distinct techniques to preprocess and encode data within a specific higher-order topological domain. This diversity complicates performance comparisons between models on different datasets. To address this issue, TopoBench implements a unifying pipeline for data preprocessing and predictive performance evaluation metrics.
Benchmarking: The described challenges collectively impede the establishment of standardized benchmarking practices within the TDL community. This work provides the first cross-domain benchmarking of TNNs across diverse datasets, adhering to a well-established and rigorous machine learning pipeline. Furthermore, TopoBench ensures the complete reproducibility of the experiments.
Democratization of TDL: The emerging nature of TDL, coupled with its reliance on advanced mathematical and computer science expertise, poses a barrier to broader adoption. TopoBench democratizes TDL by automating and modularizing the pipeline, offering a high-level interface to simplify coding, facilitating seamless integration through a modular design, and ensuring complete compatibility with the PyTorch ecosystem. It provides an accessible testbed for newcomers to experiment with topological domains, models, and datasets, fostering innovation and expanding the scope of TDL applications.
The remainder of this paper is structured as follows: Section 2 introduces key TDL concepts –with technical details in the appendix. Section 3 provides a review of related software. Section 4 details TopoBench’s modules and functionality. Section 5 demonstrates TopoBench through benchmarking experiments. Section 6 concludes with remarks and future directions.
2 Background
This section aims to build the general intuition necessary to understand TopoBench, while providing references to its formal mathematical foundations.
Figure 2: Topological Deep Learning Domains. Nodes in blue, (hyper)edges in pink, and faces in dark red. Figure adopted from Papillon et al. (2023).
Topological domains. Relational data can be represented in various forms, with graph representation being the most common framework. However, as discussed in the Introduction, graphs are limited to pairwise relations. TDL methodologies overcome this constraint by encoding higher-order relationships through combinatorial and algebraic topology concepts. Fig. 2 illustrates the standard discrete, higher-order topological spaces used to that end, which enable more complex relational representations via part-whole and set-types relations (Papillon et al., 2023); see Appendix A.1 for the formal definition of each of these topological domains.
Liftings. Since most relational data is traditionally collected in discrete domains, such as point clouds and graphs, transitioning to richer topological representations requires mappings between domains — for instance, from a graph to a simplicial complex. This process of mapping, known as lifting, enables more flexible and expressive data representations (further details in Section 4.3 and Appendix A.2).
Topological neural networks. Once the data is represented within a chosen topological domain, the TDL pipeline employs neural networks specifically designed for that domain. These models process higher-order structures, leveraging specialized inductive biases. Such networks, referred to as Topological Neural Networks (TNNs), enable learning directly from data represented through topological domains (see Appendix A.3). In general, TNNs exploit a higher-order message-passing mechanism (see Appendix A.5), which generalizes the traditional graph-based message-passing approach (see Appendix A.4), allowing for more comprehensive information propagation through higher-order structures.
@@ 3 Existing Software
Graph-based learning and GDL are supported by several software packages, including NetworkX (Hagberg et al., 2008), KarateClub (Rozemberczki et al., 2020), PyG (Fey and Lenssen, 2019), DGL (Wang et al., 2019), and Open Graph Benchmark (OGB; Hu et al., 2020, 2021). NetworkX enables computations on graphs, while KarateClub implements unsupervised learning algorithms for graph-structured data. PyG and DGL provide functionality for GDL as well as standard graph-based learning. Lastly, OGB provides a collection of graph datasets and a benchmarking framework that supports reproducible graph machine learning research; however, it does not address TDL-specific needs.
@@
Various tools also exist for higher-order domains. For hypergraphs, simplicial complexes, and other topological structures, HyperNetX (Liu et al., 2021), XGI (Landry et al., 2023), DHG (Feng et al., 2019), and TopoX (Hajij et al., 2024) each focus on different facets. HyperNetX facilitates hypergraph computations, whereas XGI supports both hypergraphs and simplicial complexes. DHG implements deep learning algorithms for graphs and hypergraphs. TopoX is a suite of three packages—TopoNetX, TopoEmbedX, and TopoModelX—that provide broader support for hypergraphs, simplicial, cellular, path, and combinatorial complexes. TopoNetX facilitates constructing and computing on these domains, including working with nodes, edges, and higher-order cells; TopoEmbedX embeds higher-order domains into Euclidean spaces, while TopoModelX implements most TNNs surveyed in Papillon et al. (2023).
Additionally, topological data analysis (TDA) libraries such as GUDHI (The GUDHI Project, 2015), giotto-tda (Tauzin et al., 2021), and scikit-tda (Nathaniel Saul, 2019) offer robust tools for topological computations, like persistent homology diagrams and topological invariant metrics. These TDA packages can provide valuable building blocks to extract topological information from data within TDL pipelines.
4 The TopoBench Library: Module Outline, Datasets and Liftings
@@
TopoBench implements a unified and flexible workflow that facilitates the addition of new datasets, data manipulation and preprocessing methods (collectively referred to as transforms), and deep learning models. This design ensures applicability across a wide range of tasks and enables a broad cross-domain comparison, currently lacking in the TDL literature. Each module within TopoBench is assigned a distinct role while maintaining a consistent input-output structure, which provides a modular interface across all topological domains. Figure 1 outlines the TopoBench modules, grouped by functionality into data, model, training, and communication components. Algorithm 1 illustrates the TopoBench execution pipeline in pseudo-code.
Algorithm 1 Execution pipeline for model training in TopoBench
@@
4.1 TopoBench Modules
Data modules. These modules manage and process data within TopoBench, including Loader, Transforms, PreProcessor, and Dataloader.
Loader. The Loader module provides an interface for downloading and storing data, built upon the widely adopted InMemoryDataset from PyG, enhancing interoperability. The project webpage offers detailed tutorials on the library, including a step-by-step guide to integrating customized data with these interfaces.
Transforms. Transforms modules are implemented as subclasses of BaseTransform (provided by PyG) and include three categories: data manipulation, topology lifting, and feature lifting. The data manipulation module enables general data transformations (e.g., adapting PyG (Fey and Lenssen, 2019) or TopoX (Hajij et al., 2024) transforms for use in TopoBench). The topology lifting and feature lifting modules handle the conversion of data from one topological domain to another (see Section 4.3). Each transform accepts a Data object as input, performs the necessary computations, and outputs the modified Data object. These composable operators can be easily customized for various tasks.
Pre-processor. The PreProcessor class applies a sequence of transforms to a dataset. It accepts a dataset object and a list of transforms, iterating over the dataset to apply each transform in turn. To avoid re-computing the same transforms repeatedly, the preprocessed dataset is saved in a dedicated folder for each transform configuration. This setup ensures that each dataset is processed only once per configuration, mitigating the potentially time-consuming nature of preprocessing large datasets. PreProcessor also generates or loads data splits according to a chosen strategy (e.g., random splits with predefined proportions, k-fold cross-validation, or fixed splits).
Dataloader. The Dataloader module provides a consistent interface for batch training across graphs, hypergraphs, simplicial complexes, cell complexes, and combinatorial complexes. By supporting mini-batching for all these domains, it helps make training more tractable on large datasets.
Model modules. The neural network modules form the core of the modeling pipeline. The encoder component maps initial data features into a latent space and applies a learnable transformation before passing the data to a TNN model –thus standardizing the input across all models. The backbone TNN can be imported from existing PyTorch libraries (e.g., TopoX or PyG), or built on a custom basis within TopoBench(see Table 8 in Appendix C). The wrapper ensures the correct input is provided to the forward pass of the backbone TNN model and collects the output in a dictionary. This design streamlines input and output handling across different topological domains, making it easier to integrate new models into TopoBench.
The readout module converts latent representations from the neural network into final predictions. The Loss module defines a loss function (from the PyTorch library, or customized), while the optimizer module configures the optimizer and scheduler. This design allows seamless use of any optimizer and scheduler from torch.optim, thereby supporting flexible and robust training. Finally, the evaluator module, built upon torchmetrics, provides metrics for both classification and regression tasks –while also allowing for tailored ones for specific datasets and tasks.
Training and communication modules. The Model class defines a training pipeline for all domains (see lines 14–19 of Algorithm 1). Inheriting from LightningModule, it requires Encoder, Wrapper, Backbone, Readout, Evaluator, Loss, and Optimizer objects as inputs. The lightning.Trainer then automates training, evaluation, and testing. Additional functionalities can be incorporated via callbacks, and users can monitor training with various loggers (e.g., wandb, tensorboard). Both are standard tools in Lightning and are referred to as communication modules in TopoBench.
4.2 Datasets
TopoBench includes a wide selection of datasets to accommodate both standard graph-based and higher-order domains. It is the first framework to enable the creation of reliable, reproducible higher-order datasets through the use of various lifting mappings. A subset of these datasets are also used in the experiments of Section 5 for demonstration purposes. See Appendix C.3 for descriptive statistics of the datasets.
Graph-based datasets. A number of well-known datasets commonly used in graph-based learning are supported. Citation networks such as Cora, Citeseer, and PubMed (Yang et al., 2016) are included, along with heterophilous datasets (where nodes connected by an edge predominantly belong to different categorical classes), such as Amazon Ratings, Roman Empire, Minesweeper, Tolokers, and Questions. The TU datasets, including MUTAG, PROTEINS, NCI1, NCI109, IMDB-BIN, IMDB-MUL, and REDDIT (Morris et al., 2020), are also integrated, as are molecule datasets like ZINC (Gómez-Bombarelli et al., 2018) and AQSOL (Dwivedi et al., 2023). Furthermore, TopoBench supports the US County Demographics dataset (Jia and Benson, 2020), illustrating its adaptability to various graph structures.
Datasets with higher-order interactions. Several datasets with higher-order interactions are included in TopoBench, showcasing its capabilities to handle data supported on hypergraphs, simplicial complexes, and other topological domains. The Mantra dataset (Ballester et al., 2024) is part of TopoBench, offering over 43,000 two-dimensional and 249,000 three-dimensional triangulations of surfaces and manifolds, which can be used, for example, as features on a simplicial complex. In addition, the widely used AllSet hypergraph datasets (Chien et al., 2021)—Cora-Cocitation, Citeseer-Cocitation, PubMed-Cocitation, Cora-Coauthorship, and DBLP-Coauthorship—are integrated, following the same preprocessing as HyperGCN (Yadati et al., 2019). These hypergraphs group documents co-authored or co-cited together into single hyperedges. Collectively, these examples illustrate how TopoBench supports data beyond traditional graph pairwise interactions.
Compatibility and custom datasets. To simplify dataset integration, TopoBench provides convenient wrappers that build on PyG loaders (e.g., TUDatasets, Planetoid, ZINC). In many cases, these wrappers enable straightforward use of any graph dataset already supported by PyTorch Geometric, as well as newly introduced datasets such as Mantra, the hypergraph citation networks, and Human3.6m. Support for custom datasets is facilitated by a simple interface with two key methods: download(), for fetching or extracting raw files, and process(), for converting the data into the desired relational structure (graph, hypergraph, simplicial or cell complex). Code examples and tutorials provided in TopoBench illustrate the TopoBench interface for loading custom user-defined datasets [^2] https://github.com/geometric-intelligence/TopoBench/blob/main/tutorials/tutorial_add_custom_dataset.ipynb. This approach guarantees users can easily extend TopoBench to any dataset of interest, thus maintaining the library’s modular and extensible design.